OpenStudy (anonymous):

Identify the y-intercept in the following equation y = 8x + 4.

OpenStudy (anonymous):

make y=o

OpenStudy (anonymous):

b = 4 which is the y intercept

OpenStudy (anonymous):

and m = 8

OpenStudy (anonymous):

0=8x+4

OpenStudy (anonymous):

4/8=x?

OpenStudy (jjuden):

Slope=8 Y intercept=4

OpenStudy (anonymous):

Okay, can you help me with another?

OpenStudy (jjuden):

yea

OpenStudy (anonymous):

Which of the following is the equation of a line with a slope of -2 that passes through (-1, 3)? I got to: y-3=-2(x+1)

OpenStudy (anonymous):

then it's y=1(x+1)... Then after that I'm not sure

OpenStudy (jjuden):

ok hold on

OpenStudy (jjuden):

ok wat other equations are there for answers cause yours is wrong

OpenStudy (anonymous):

a. y = -2x+1 b. y = -2x+5 c. y = -2x+3 d. y = 2x+1

OpenStudy (jjuden):

the answer is A

OpenStudy (anonymous):

Thanks, one more?

OpenStudy (jjuden):

ok

OpenStudy (anonymous):

Choose the property that justifies: -7x = -7x

OpenStudy (anonymous):

a. Distributive Property for Real Numbers b. Reflexive Property of Equality c. Substitution Property of Equality d. Symmetric Property of Equality e. Transitive Property of Equality

OpenStudy (jjuden):

B.Reflexive Property of Equality It basically states that an expression is equal to itself c = c

OpenStudy (anonymous):

thanks!

OpenStudy (jjuden):

welcome :)

OpenStudy (jjuden):

if you ever need to graph something you are stuck on use this website http://www.quickmath.com/webMathematica3/quickmath/graphs/equations/basic.jsp#c=graph_stepsgraphequation i use this when im stuck